OpenStudy (anonymous):

Which of the following is equivalent to the statement below? If 3x + 1 = 10, then x ≠ 2.

OpenStudy (anonymous):

.A) If x ≠ 2, then 3x + 1 = 10. .B) If 3x + 1 = 10, then x = 2. .C) If x = 2, then 3x + 1 = 10. .D) If x = 2, then 3x + 1 ≠ 10.

OpenStudy (anonymous):

Put 2 into your formula, what is the answer?

OpenStudy (anonymous):

equivalent statement is the contrapositive

OpenStudy (anonymous):

contrapositive of if P then Q is if not Q then not P

OpenStudy (anonymous):

Or listen to him. :P

OpenStudy (anonymous):

If 3x + 1 = 10, then x ≠ 2 P: \(3x+1 = 10\) Q: \(x \neq 2\)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\(\lnot Q: x=2\) \(\lnot P: 3x+1\neq 10\)

OpenStudy (anonymous):

you have a statement of the form "if P then Q"

OpenStudy (anonymous):

If x ≠ 2, then 3x + 1 = 10.?

OpenStudy (anonymous):

where P is \( 3x+1=10\) and Q is \(x\neq 2\)

OpenStudy (anonymous):

no first you have to negate each of these statements

OpenStudy (anonymous):

the negation of \(3x+1=10\) is \(3x+1\neq 10\) and the negation of \(x\neq 2\) is \(x=2\)

OpenStudy (anonymous):

If x = 2, then 3x + 1 = 10.?

OpenStudy (anonymous):

so to recap slowly P: \(3x+1=10\) not P: \(3x+1\neq 10\) Q: \(x\neq 2\) not Q: \(x=2\)

OpenStudy (anonymous):

huh? im confusseed

OpenStudy (anonymous):

you need "if not Q then not P"

OpenStudy (anonymous):

and since not Q is \(x=2\) then the contrapostive is "if \(x=2\) then \(3x+1\neq 10\) which is certainly a true statement

OpenStudy (anonymous):

a statement "if P then Q" is equivalent to the contrapositive statement "if not Q then not P"

OpenStudy (anonymous):

you following or still a mystery?

OpenStudy (anonymous):

mystery

OpenStudy (anonymous):

ok we start with the original statement if \(3x+1=10\) then \(x\neq 2\)

OpenStudy (anonymous):

it is of the form if P then Q

OpenStudy (anonymous):

what is P?

OpenStudy (anonymous):

3x + 1 = 10

OpenStudy (anonymous):

ok good now what is "not P"

OpenStudy (anonymous):

if that is not obvious, let me know and i will tell you

OpenStudy (anonymous):

x ≠ 2.

OpenStudy (anonymous):

If 3x + 1 = 10, then x = 2.